A 1,386 kilogram truck pulling a 820 kilogram trailer exerts a 2,363 Newton force on the road and produces an acceleration of 0.6 meters/second 2. There is clearly some resistance force, such as friction and air resistance. Assume that 70 % of this resistance force is exerted on the trailer. What is the horizontal force in the hitch between the truck and the trailer measured in Newtons? Hint: The correct answer is a positive number.
Answer #1

Total resistance = f

F = 2363 N

total mass = 1386 + 820 = 2206 kg

acceleration = 0.6

using newtons law

F - f = ma

f = F - ma = 2363 - 2206*0.6 = 1039.4 N

Now for trailer only

let hitch force = H

resistance = 0.7*f = 727.58 N

again using newtons law

H - 727.58 = 820*0.6

H = 1219.58 N

Please upvote if you like the answer!!